# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

### v3

@@ Abstract
Machine learning models traditionally assume that training and test data are independently and identically distributed. However, in real-world applications, the test distribution often differs from training. This problem, known as out-of-distribution (OOD) generalization, challenges conventional models. Invariant Risk Minimization (IRM) emerges as a solution that aims to identify invariant features across different environments to enhance OOD robustness. However, IRM’s complexity, particularly its bi-level optimization, has led to the development of various approximate methods. Our study investigates these approximate IRM techniques, employing the Expected Calibration Error (ECE) as a key metric to measure the extent to which the model has acquired invariant features. ECE, which measures the reliability of model prediction, serves as an indicator of whether models effectively capture environment-invariant features. Through a comparative analysis of datasets with distributional shifts, we observe that Information Bottleneck-based IRM, which condenses representational information, achieves a balance in improving ECE while preserving accuracy relatively. This finding is pivotal, demonstrating a feasible path to maintaining robustness without compromising accuracy. Nonetheless, our experiments also caution against over-regularization, which can diminish accuracy. This underscores the necessity for a systematic approach in evaluating OOD generalization metrics, which goes beyond mere accuracy to address the nuanced interplay between accuracy and calibration. Our code is available at https://github.com/katoro8989/IRM_Variants_Calibration
@@ 1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Vapnik, 1991) and calibration (Guo et al., 2017a) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge. Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify constant features across different environments, facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020; Zhang et al., 2023a). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution (ID) and OOD accuracy. This implies that models trained by IRM variants failed to obtain invariant features since they sacrificed ID accuracy to achieve better OOD accuracy. From another perspective of IRM, it can also be conceptualized as a specialized form of multi-domain calibration (Wald et al., 2021a). Our study intersects with these findings, examining how IRM variants mimic the original model in calibration. We posit that effective implementation of IRM should lead to reduced loss and a zero expected calibration error (ECE) (Wald et al., 2021a). Through comparative experimentation, we assess various IRM variants in environments marked by distributional shifts, such as diversity and correlation shifts (Ye et al., 2022). Our evaluation focuses not only on accuracy but also on ECE. The results highlight notable disparities in ECE among different methodologies, even when OOD accuracy levels are similar. This suggests that a sole reliance on OOD accuracy for method development might be inadequate. Our observations include a range of calibration behaviors across methods, from overconfidence to underestimation.
@@
  • We study discrepancies between IRM formulation and its implemented variants using ECE as a key metric to measure how much the model learned environmental invariant features.
  • We reveal that IRM variants with strong regularization typically achieve lower ECE (Figure 3 and Figure 6).
  • We observed that Information Bottleneck-based IRM (IB-IRM) lowers ECE across the environments (Figure 5(c)), aligning it more closely with the original objectives of IRM while maintaining OOD accuracy relatively (Figure 2 and Table 2).
  • In contrast, for IRMv1, we identify a trade-off between ECE and accuracy, thereby underscoring the compromises necessary for achieving robust generalization and accurate calibration (Figure 4).
Our study, using ECE as a metric, illuminates the gap between the original formulation of IRM and its variants. We recognize a notable trade-off between ECE and accuracy, highlighting the complexities and necessary compromises in striving for robust generalization and precise calibration.
@@ 2.1 OOD Generalization
$\min_{f:\mathcal{X}\to\mathcal{Y}}\sum_{e\in E_{train}}R^{e}(f)$ (1)
@@ 2.2 Calibration
The alignment of a model’s predictive confidence with its actual accuracy is extremely important in practical scenarios such as medical image diagnosis and autonomous driving. For example, if there are 100 instances where a model predicts with 80% confidence, then 80 of those predictions should be correct. However, in recent deep neural networks, issues such as increased model capacity have led to problems with miscalibration, and many methods have been proposed to address this calibration issue (Guo et al., 2017a). The correct calibration in all environments is defined for any predictive probability $\alpha$ of $f$ as follows:
@@
Furthermore, the quality of uncertainty calibration is often quantified as ECE and is commonly used. It is defined as follows:
@@
$Acc(B_{m})=$ $\frac{1}{\left|B_{m}\right|}\sum_{b\in B_{m}}\mathbf{1}\left(\hat{y}_{b}=y_{b}\right)$
$Conf(B_{m})=$ $\frac{1}{\left|B_{m}\right|}\sum_{b\in B_{m}}\hat{p}_{b}$
@@ 2.3 OOD Generalization and Calibration
It has been pointed out that OOD generalization and calibration are interrelated. Notably, eq. 3 can be a sufficient condition for eq. 2, theoretically demonstrating that calibration across all environments leads to OOD generalization. Wald et al. (2021b) capitalize on this observation and propose CLOvE, a method that calibrates models across multiple environments, thereby enabling generalization to unseen data distributions.
Additionally, it has been empirically demonstrated that there is a correlation between a model’s accuracy in OOD scenarios and its calibration performance. It has been demonstrated that there exists a correlation between a model’s accuracy and its ECE under dataset shift scenarios. Furthermore, techniques such as temperature scaling (Guo et al., 2017b), originally proposed for calibration purposes, have been shown to be effective for OOD generalization (Ovadia et al., 2019; Naganuma & Hataya, 2023).
In this paper, we utilize this fact and hypothesize that models with high calibration performance will likely achieve better OOD generalization. We evaluate the approximation methods of Invariant Risk Minimization (IRM), introduced in Section 4, based on the model’s calibration performance to assess their OOD generalization capabilities.
@@ 4.3 Pareto IRM(PAIR)
$\mathcal{L}_{ERM}$ $=\sum_{e\in E_{train}}R^{e}(\Phi^{\prime})$
$\mathcal{L}_{IRMv1}$ $=\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2}$
$\mathcal{L}_{VREx}$ $=Var(\{R^{e}(\Phi^{\prime})\}_{e\in E_{train}})$
@@ 4.5 Bayesian IRM(BIRM)
ERM IRMv1 BIRM IB-IRM
IRM Game PAIR 100% Confidence 50% Confidence
Figure 2: Comparison of OOD accuracy (horizontal axis) and OOD ECE (vertical axis) in CMNIST. The red solid line represents the theoretical values when all the model’s predictive probabilities are 100%, and the blue solid line represents the case when all predictive probabilities are 50%. IRMv1, IB-IRM, and BIRM, which achieve relatively high inference accuracy, are distributed near the blue line. This indicates that these models mitigate reliance on non-invariant features by making predictions more ambiguous.
@@ 5 Experiments
Although the approximation methods in the previous chapter have achieved some success in terms of OOD accuracy on OOD datasets (Arjovsky et al., 2020; Ahuja et al., 2022; Chen et al., 2022; Ahuja et al., 2020; Lin et al., 2022), it is not known how invariant features they actually acquire and how robust they are to unknown environments. Therefore, we evaluated them by comparing their calibration performance with ECE.
@@ 5.2.2 Results
Figure 3: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 eq. 6. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in a lower ECE for the model in the OOD environment.
Figure 2 presents a scatter plot showing the relationship between OOD accuracy and ECE for multiple IRM variants, each trained with multiple hyperparameters. The red solid line represents the theoretical value when all model’s predictive probabilities are 100%, indicating very high confidence. In contrast, the blue solid line represents the case when all predictive probabilities are 50%, indicating very low and ambiguous confidence. The methods with relatively high accuracy (IRMv1(green), IB-IRM(light blue), BIRM(red)) align with the blue solid line, showing that they can overcome excessive fitting to non-invariant features by making predictions more ambiguous. On the other hand, ERM(blue) and IRM Game(purple) tend to be distributed near the red solid line, indicating predictions with relatively high confidence. PAIR(yellow) is distributed between the two solid lines, demonstrating intermediate characteristics between the two aforementioned groups.
Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of OOD generalization actually affects the ECE. It specifically compares different values of $\lambda$ in IRMv1’s formulation eq. 6, demonstrating that as the impact of IRMv1’s penalty increases, the ECE decreases, suggesting better calibration.
@@ 5.3.2 Results
Figure 4: The relationship between OOD accuracy and ECE varies with the $\lambda$ parameter of IRMv1 in the RMNIST dataset. As the color shifts from blue to green, $\lambda$ increases, indicating stronger regularization in the model. Points closer to blue, situated on the middle right of the figure, correspond to intermediate ECE and high accuracy. In contrast, points nearer to green, located towards the lower left, indicate better ECE but lower accuracy. This suggests that in more practical datasets, such as those with diversity shifts, stronger regularization may improve ECE, but this often comes at the cost of a significant decrease in accuracy.
Considering that a model’s OOD accuracy affects its calibration (Ovadia et al., 2019), a comparative evaluation of ECE combined with OOD accuracy was conducted for a fairer comparison of ECE. A threshold was set for validation accuracy in the training environment, and upon reaching this threshold, early stopping of training was implemented to create conditions of comparable OOD accuracy.
The average results from three different random seeds under the aforementioned conditions are presented in Table 2, where although all methods achieved a similar average OOD accuracy, notable differences were observed in OOD ECE. Specifically, the ECE of IB-IRM was, on average, lower than that of any other method. Additionally, Figure 5(c) visualizes the relationship between ECE in the training and test environments, with the red solid line representing cases where ECE is equal in both environments. Ideally, models should be distributed near this red line from the perspective of cross-environmental calibration of uncertainty. Specifically, in Figure 5(a) and Figure 5(c), IB-IRM is distributed near the red solid line compared to other methods, suggesting that it does not overfit to the training environment and exhibits more ideal feature learning.
@@
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ERM IRMv1 BIRM IB-IRM IRM Game PAIR ID-ECE = OOD-ECE
It was also suggested that regularization in models such as IRMv1 and IBIRM, while reducing OOD ECE in cases of diversity shift, can adversely affect accuracy. Figure 4 illustrates the relationship between OOD accuracy and ECE in RMNIST when varying the magnitude of $\lambda$ in IRMv1. As regularization strengthens, the model’s ECE improves; however, this is accompanied by a drastic deterioration in accuracy, indicating a trade-off between the two.
@@
OOD ECE RMNIST 5.27
$\pm$1.11
3.98
$\pm$0.69
1.58 $\pm$0.44 6.93
$\pm$0.56
7.17
$\pm$0.98
7.88
$\pm$0.79
PACS 11.99 $\pm$1.86 13.35
$\pm$1.45
12.85
$\pm$0.60
13.73
$\pm$1.93
12.50
$\pm$1.45
12.46
$\pm$0.95
VLCS 12.53
@@
11.90
$\pm$1.26
7.84 $\pm$1.55 10.57
$\pm$1.81
8.70
$\pm$0.93
10.34
$\pm$1.11
Avg. 9.93
$\pm$1.46
9.74
$\pm$1.13
7.42 $\pm$0.86 10.41
$\pm$1.43
9.46
$\pm$1.12
10.23
$\pm$0.95
Table 2: Comparison of OOD accuracy and ECE across RMNIST, PACS, and VLCS datasets. Following hyperparameter tuning for each method, measurements were taken across three different random seeds, presenting the mean and variance for each metric. It was observed that on average, IB-IRM exhibits a lower ECE while maintaining comparable inference performance.

### v4

@@ Abstract
Machine learning models traditionally assume that training and test data are independently and identically distributed. However, in real-world applications, the test distribution often differs from training. This problem, known as out-of-distribution (OOD) generalization, challenges conventional models. Invariant Risk Minimization (IRM) emerges as a solution that aims to identify invariant features across different environments to enhance OOD robustness. However, IRM’s complexity, particularly its bi-level optimization, has led to the development of various approximate methods. Our study investigates these approximate IRM techniques, using the consistency and variance of calibration across environments as metrics to measure the invariance aimed for by IRM. Calibration, which measures the reliability of model prediction, serves as an indicator of whether models effectively capture environment-invariant features by showing how uniformly over-confident the model remains across varied environments. Through a comparative analysis of datasets with distributional shifts, we observe that Information Bottleneck-based IRM achieves consistent calibration across different environments. This observation suggests that information compression techniques, such as Information Bottleneck are potentially effective in achieving model invariance. Furthermore, our empirical evidence indicates that models exhibiting consistent calibration across environments are also well-calibrated. This demonstrates that invariance and cross-environment calibration are empirically equivalent. Additionally, we underscore the necessity for a systematic approach to evaluating OOD generalization. This approach should move beyond traditional metrics, such as accuracy and F1 scores, which fail to account for the model’s degree of over-confidence, and instead focus on the nuanced interplay between accuracy, calibration, and model invariance. Our code is available at https://github.com/katoro8989/IRM_Variants_Calibration
@@ 1 Introduction
In the evolving landscape of machine learning, the importance of out-of-distribution (OOD) generalization (Vapnik, 1991) and calibration (Guo et al., 2017) cannot be understated, especially in real-world applications and critical scenarios. Traditional machine learning approaches, grounded in the assumption that data are independently and identically distributed (IID), often struggle to cope with OOD scenarios. Additionally, a notable trend in contemporary machine learning models is their tendency towards overconfidence, which undermines the reliability of their confidence estimations. This issue is primarily recognized as a calibration challenge. Responding to these limitations, there has been a surge in research focusing on methodologies like Invariant Risk Minimization (IRM) (Arjovsky et al., 2020). IRM is designed to identify constant features across different environments, facilitating more robust generalization in the face of environmental variations. However, the computational demands of IRM have led to the exploration of more feasible, approximate methods. Despite the development of several IRM variants, their inability to consistently surpass the performance of finely-tuned Empirical Risk Minimization (ERM) models has been observed (Gulrajani & Lopez-Paz, 2020; Zhang et al., 2023a). This performance gap raises critical questions about the inherent deficiencies of these IRM variants. As we showed in Figure 1, even by leveraging IRM variants, there is a trade-off between in-distribution (ID) and OOD accuracy. This implies that models trained by IRM variants failed to obtain invariant features since they sacrificed ID accuracy to achieve better OOD accuracy. The invariance aimed for by IRM can, from another perspective, be conceptualized as the invariance of calibration degree, meaning that the model remains equally over-confident across all environments (see Section 2.3 for details). Our study intersects with this insight by evaluating the invariance of calibration through the variance of calibration metrics across environments, examining the extent to which IRM variants mimic the original model in terms of invariance. We hypothesize that an effective implementation of IRM should result in consistent calibration across all environments, indicated by minimal variance in calibration. The results reveal significant disparities in how calibration performance behaves across different environments depending on the IRM methodology used, even when OOD accuracy levels are similar. This suggests that relying solely on OOD accuracy for method development may be inadequate.
@@
  • We study discrepancies between IRM formulation and its implemented variants using the calibration metrics as a key metric to measure how much the model learned environmental invariant features.
  • We observed that Information Bottleneck-based IRM (IB-IRM) improves the calibration performance across the environments (Figure 4(c) and Figure 2(c)).
  • We demonstrate that the trend of IB-IRM aligns more closely with the original objectives of IRM since its ECE variants across all environments are the smallest (Table 2).
@@ 2.1 OOD Generalization
$\vspace{-3mm}\min_{f:\mathcal{X}\to\mathcal{Y}}\sum_{e\in E_{train}}R^{e}(f)\vspace{-1mm}$ (1)
@@ 2.2 Calibration
The alignment of a model’s predictive confidence with its actual accuracy is extremely important in practical scenarios such as medical image diagnosis and autonomous driving. For example, if there are 100 instances where a model predicts with 80% confidence, then 80 of those predictions should be correct. However, in recent deep neural networks, issues such as increased model capacity have led to problems with miscalibration, and many methods have been proposed to address this calibration issue (Guo et al., 2017). The correct calibration in all environments is defined for any predictive probability $\alpha$ of $f$ as follows:
@@
Expected Calibration Error
The quality of uncertainty calibration is often quantified as ECE (Naeini et al., 2015) and is commonly used. It is defined as follows:
@@
$Acc(B_{m})=\frac{1}{\left|B_{m}\right|}\sum_{b\in B_{m}}\mathbf{1}\left(\hat{y}_{b}=y_{b}\right),\ Conf(B_{m})=\frac{1}{\left|B_{m}\right|}\sum_{b\in B_{m}}\hat{p}_{b}$
@@
Adaptive Calibration Error
Nixon et al. (2019) claimed that ECE is not perfect for measuring model calibration performance in some cases. Specifically, ECE calculates the discrepancy between predicted probabilities and actual accuracy for fixed bins based on predicted probabilities. If the model’s predictions exhibit sharpness, this can result in a data imbalance within each bin. This trend of sharpness has been particularly noted in recent large-scale neural networks. The imbalance in the number of data points within bins implies a bias in the quality of calibration estimation, potentially leading to inaccurate assessments. To address the aforementioned issue, the Adaptive Calibration Error (ACE) has been proposed (Nixon et al., 2019). Unlike ECE, ACE divides the bins such that each bin contains the same number of samples, rather than being based on predicted probabilities. The formulation is as follows:
$ACE=\frac{1}{KM}\sum_{k=1}^{K}\sum_{m=1}^{M}|Acc(B_{m,k})-Conf(B_{m,k})|$ (5)
where $K$ denotes the number of classes. The predicted probabilities corresponding to class $k$ for each data point are sorted and divided equally into $M$ bins, with the $m$-th bin denoted as $B_{m,k}$. In other words, each bin contains $\lfloor N/M\rfloor$ samples.
However, it should be noted that, in our scope of experiments, there is no significant difference between ECE and ACE results.
Negative Log-Likelihood
Additionally, the Negative Log-Likelihood (NLL) is commonly used as an indirect metric for measuring calibration performance (Minderer et al., 2021). The NLL is expressed by the following equation, and the closer the model’s confidence in the correct label is to 1, the more favorable it is.
$NLL=-\sum_{n=1}^{N}\log\hat{p}(y_{n}|x_{n},\theta)$ (6)
Therefore, even if the accuracy is 100%, improving the model’s confidence can still enhance the NLL. If the model is overfitting on the NLL in the training domain and becoming overconfident, it is more likely to make incorrect predictions with overconfidence in the test domain. This leads to a higher NLL, indicating poor calibration performance (Mukhoti et al., 2020).
@@ 2.3 OOD Generalization and Calibration
Some empirical study (Naganuma & Hataya, 2023) reveals the correlation between OOD generalization and calibration. Theoretically, eq. 3 can be a sufficient condition for eq. 2, demonstrating that calibration across all environments leads to OOD generalization. Wald et al. (2021) capitalize on this observation and propose CLOvE, a method that calibrates models across multiple environments, thereby enabling generalization to unseen data distributions.
Why do we use calibration to evaluate IRM variants?
Since eq. 3 is a sufficient condition for eq. 2 and not a necessary one, it is possible for a model to achieve OOD while making overconfident predictions. Therefore, calibration performance does not necessarily perfectly explain OOD performance.
However, comparing the degree of calibration across different environments allows for measuring a model’s invariance. The invariance condition is defined by the equality of the conditional probability $E[Y_{e}|f(X_{e})]$ across all environments, which can be restated as the consistency of a model’s miscalibration across all environments. Thus, evaluating the variance of a model’s calibration metric across environments provides a means to assess its invariance. For example, if a model is consistently over-confident by 0.1 across all environments, meaning that $\mathbb{E}[Y^{e}|f(X^{e})=\alpha]=\min(1.0,\alpha+0.1)\ \forall e\in E_{all}$ holds, then the model satisfies invariance.
How do we approximate invariance condition by using ECE?
Each term of invariance condition eq. 2, can be approximated by the conditional accuracy, $\mathbb{E}[Y_{e}|f(X_{e})]\approx Acc({B^{e}}_{m})$. Then, Equation 2 could be approximated as $Acc({B^{e}}_{m})=Acc({B^{e^{\prime}}}_{m})\forall e,e^{\prime}\in E_{all}$. And this equation can be easily transformed into $Conf({B^{e}}_{m})-Acc({B^{e}}_{m})=Conf({B^{e^{\prime}}}_{m})-Acc({B^{e^{\prime}}}_{m})\forall e,e^{\prime}\in E_{all}$. Given that recent models are empirically overconfident (Minderer et al., 2021), we assume $Conf(B^{e}_{m})>Acc(B^{e}_{m})$, so the equation $Gap(B^{e}_{m})=Conf(B^{e}_{m})-Acc(B^{e}_{m})=|Acc(B^{e}_{m})-Conf(B^{e}_{m})|$ holds true. Therefore, Equation 2 can be approximated as $ECE_{e}=ECE_{e^{\prime}}\forall e,e^{\prime}\in E_{all}$, and the consistency of ECE across environments can be used to approximate the model’s invariance.
Furthermore, our empirical results indicate that cases of equivalent calibration failure between ID and OOD are rare. When calibration is consistent across domains, it is typically in well-calibrated cases. Hence, empirical evidence suggests an equivalence between the invariance condition and calibration across all environments.
In this paper, we utilize this fact that comparing calibration performance across different environments will provide a better assessment of OOD generalization capabilities and evaluate the approximation methods of IRM, introduced in Section 4.
@@ 4.3 Pareto IRM(PAIR)
$\mathcal{L}_{ERM}=\sum_{e\in E_{train}}R^{e}(\Phi^{\prime}),\ \ \mathcal{L}_{IRMv1}=\parallel\nabla_{\omega^{\prime}\mid\omega^{\prime}=1.0}R^{e}(\omega^{\prime}\cdot\Phi^{\prime})\parallel^{2},\ \ \mathcal{L}_{VREx}=Var(\{R^{e}(\Phi^{\prime})\}_{e\in E_{train}})$
@@ 4.5 Bayesian IRM(BIRM)
(a) ECE vs ECE
(b) ACE vs ACE
(c) NLL vs NLL
Figure 2: ECE, ACE, and NLL evaluation on the CMNIST. The X-axis indicates the evaluation of ID sets, and the Y-axis indicates the evaluation of OOD sets. The red solid line represents the case when the metric values are equal for both ID and OOD, indicating that the model has achieved the same level of calibration performance in both domains. IRMv1, IB-IRM, BIRM, and PAIR are distributed relatively close to the red solid line, indicating that they achieve consistent calibration across environments. Those data points that show better OOD calibration performance for each IRM variant tend to be closer to the red line.
{mdframed}
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ERM IRMv1 BIRM IB-IRM IRM Game PAIR ID = OOD
@@ 5 Experiments
Although the approximation methods in the previous chapter have achieved some success in terms of OOD accuracy on OOD datasets (Arjovsky et al., 2020; Ahuja et al., 2022; Chen et al., 2022; Ahuja et al., 2020; Lin et al., 2022), it is not known how invariant features they actually acquire and how robust they are to unknown environments. Therefore, we evaluated them by comparing their calibration performance with ECE, ACE, and NLL. Since we aim to investigate OOD performance through invariance condition eq. 2, we do not use some metrics that do not incorporate the model’s conditional probability, such as the F1 score.
@@ 5.2.2 Results
Figure 3: Comparison of the model’s OOD calibration performance in CMNIST, with various values of $\lambda$ in the formulation of IRMv1 eq. 8. It was observed that increasing $\lambda$, thereby increasing the regularization penalty in IRMv1, results in improved calibration performance for the model in the OOD environment.
Figure 2(c) presents a scatter plot showing the relationship between evaluation on ID and that on OOD for multiple IRM variants, each trained with multiple hyperparameters. The red solid line represents the case when the metric values are equal for both ID and OOD, indicating that the model has achieved the same level of calibration performance in both domains. IRMv1 (green), IB-IRM (light blue), BIRM (red), and PAIR (yellow) are distributed near the red solid line. Additionally, they exhibit better calibration performance in both domains. On the other hand, ERM (blue) and IRM Game (purple) are distributed above the red solid line, showing worse calibration performance in OOD. This suggests that they may be overfitting to the training domain, resulting in overconfidence in the test domain.
In some cases, IRM Game shows poor calibration performance but still appears near the red solid line. However, it does not exhibit stable distribution. This instability is likely due to the presence of separate classifiers $\omega^{e}$ fitted for each environment, which may lead to overfitting within each environment, making it difficult to balance across them.
Empirically, cases, where calibration performance is consistent between ID and OOD, are limited to those where both domains exhibit better calibration performance. Specifically, as shown in Figure 2(c), instances distributed along the red solid line are confined to points with low NLL in both domains. This trend is also observed in ECE (Figure 2(a)) and ACE (Figure 2(b)), suggesting that, empirically, cross-environment calibration is a necessary condition for invariance.
Moreover, Figure 3 visualizes how the penalty introduced in IRMv1 for the purpose of OOD generalization actually affects the calibration performance. It specifically compares different values of $\lambda$ in IRMv1’s formulation eq. 8, demonstrating that as the impact of IRMv1’s penalty increases, the calibration performance improves. It was also found that the information bottleneck penalty in IB-IRM is effective in improving calibration. Detailed results are presented in Figure 7 and Figure 8 in Section B.2.
@@ 5.3.2 Results
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ERM IRMv1 BIRM IB-IRM IRM Game PAIR ID = OOD
Considering that a model’s OOD accuracy affects its calibration (Ovadia et al., 2019), we compared the calibration performance with OOD accuracy for a fairer comparison of the calibration metrics. A threshold was set for validation accuracy in the training environment, and upon reaching this threshold, early stopping of training was implemented to create conditions of comparable OOD accuracy. The threshold for early stopping was selected using grid search to minimize the variance in test accuracy across different methods.
Table 2 shows the variance in calibration performance across all environments for each method on different datasets. Despite similar OOD accuracy, the calibration performance behavior varies across methods. Notably, IB-IRM consistently achieves a lower variance in most cases, indicating that it maintains similar calibration across all environments and achieves invariance. In Figure 4(c), it is also evident that IB-IRM achieves similar calibration performance for both ID and OOD. Additional ACE and NLL results are presented in Section B.1.
Similar to the Correlation Shift, it was observed that when calibration performance is consistent across both domains, there is a strong tendency to achieve good calibration performance. Detailed explanations along with more experimental results are provided in Section B.1.
@@
ECE Variance RMNIST 2.21
$\pm$0.97
3.61
$\pm$1.73
0.17 $\pm$0.03 1.85
$\pm$0.49
4.37
$\pm$1.79
11.17
$\pm$3.04
PACS 19.86
$\pm$17.38
23.62
$\pm$1.77
22.10
$\pm$6.39
45.78
$\pm$14.01
16.73
$\pm$6.05
6.19 $\pm$7.49
VLCS 17.97
$\pm$2.51
65.99
$\pm$5.82
5.80 $\pm$3.41 12.84
$\pm$5.57
6.49
$\pm$2.46
33.50
$\pm$19.74
Avg. 13.35
$\pm$6.95
31.07
$\pm$3.11
9.36
$\pm$3.28
20.16
$\pm$6.69
9.22 $\pm$27.59 16.95
$\pm$10.09
ACE Variance RMNIST 2.31
$\pm$0.87
3.68
$\pm$1.79
0.18 $\pm$0.08 1.85
$\pm$0.48
4.37
$\pm$2.04
11.65
$\pm$2.96
PACS 21.15
$\pm$18.69
25.33
@@
20.24
$\pm$7.27
47.31
$\pm$13.65
18.94
$\pm$5.86
6.29 $\pm$6.73
VLCS 20.20
$\pm$2.23
67.65
$\pm$6.06
6.29 $\pm$3.59 12.29
$\pm$4.67
6.98
$\pm$1.88
33.90
$\pm$19.90
Avg. 14.55
$\pm$7.26
32.22
$\pm$3.09
8.90 $\pm$3.65 20.48
$\pm$6.27
10.10
$\pm$3.26
17.28
$\pm$9.86
NLL Variance RMNIST 0.01 $\pm$0.00 0.02
$\pm$0.01
0.02
$\pm$0.00
0.02
$\pm$0.00
0.03
$\pm$0.02
0.05
$\pm$0.01
PACS 0.15
$\pm$0.13
0.12
$\pm$0.04
0.09
$\pm$0.01
0.23
$\pm$0.06
0.12
$\pm$0.01
0.05 $\pm$0.01
VLCS 0.12
$\pm$0.01
0.32
$\pm$0.01
0.10 $\pm$0.01 0.11
$\pm$0.01
0.10 $\pm$0.01 0.19
$\pm$0.03
Avg. 0.09
$\pm$0.05
0.15
$\pm$0.02
0.07 $\pm$0.01 0.12
$\pm$0.02
0.08
$\pm$0.01
0.10
$\pm$0.02
Table 2: Comparison of calibration performance consistency across environments for RMNIST, PACS, and VLCS. This analysis presents the variance in ECE, ACE, and NLL across environments, where smaller variances indicate more consistent calibration across environments. Each model was trained with hyperparameter tuning using three different seeds, and the average variance across environments was calculated. The results show that IB-IRM achieved the smallest average variance, indicating that it consistently maintains calibration across different environments.